# Practical Region-level Attack against Segment Anything Models

### v1

@@ Abstract
Segment Anything Models (SAM) have made significant advancements in image segmentation, allowing users to segment target portions of an image with a single click (i.e., user prompt). Given its broad applications, the robustness of SAM against adversarial attacks is a critical concern. While recent works have explored adversarial attacks against a pre-defined prompt/click, their threat model is not yet realistic: (1) they often assume the user-click position is known to the attacker (point-based attack), and (2) they often operate under a white-box setting with limited transferability. In this paper, we propose a more practical region-level attack where attackers do not need to know the precise user prompt. The attack remains effective as the user clicks on any point on the target object in the image, hiding the object from SAM. Also, by adapting a spectrum transformation method, we make the attack more transferable under a black-box setting. Both control experiments and testing against real-world SAM services confirm its effectiveness.
{strip}
@@ 1 Introduction
In this paper, we introduce a region-level attack to explore a more practical threat model (see Practical Region-level Attack against Segment Anything Models). The attacker’s goal is to conceal the object within an attacker-specified region from SAM’s segmentation. In this case, the attacker does not need to know the precise point of the click of the user—no matter which point in the region is clicked by the user, the object cannot be accurately segmented by SAM. In addition, we investigate to improve the transferability of the attack such that it can operate under a black-box setting. As shown in Practical Region-level Attack against Segment Anything Models, after applying the adversarial perturbation to the image, clicking on any point in the yellow box (attacker-specified region) will no longer separate the corgi from the rest of the image.
@@ 3 Preliminary
[…]
@@ 4.1 Threat Movel and Attack Overview
[…]
@@
[…]

### v2

@@ Abstract
Segment Anything Models (SAM) have made significant advancements in image segmentation, allowing users to segment target portions of an image with a single click (i.e., user prompt). Given its broad applications, the robustness of SAM against adversarial attacks is a critical concern. While recent works have explored adversarial attacks against a pre-defined prompt/click, their threat model is not yet realistic: (1) they often assume the user-click position is known to the attacker (point-based attack), and (2) they often operate under a white-box setting with limited transferability. In this paper, we propose a more practical region-level attack where attackers do not need to know the precise user prompt. The attack remains effective as the user clicks on any point on the target object in the image, hiding the object from SAM. Also, by adapting a spectrum transformation method, we make the attack more transferable under a black-box setting. Both control experiments and testing against real-world SAM services confirm its effectiveness. Code is released at https://github.com/ShenYifanS/S-RA_T-RA.
{strip}[Uncaptioned image]
@@ 1 Introduction
In this paper, we introduce a region-level attack to explore a more practical threat model (see Fig. 1). The attacker’s goal is to conceal the object within an attacker-specified region from SAM’s segmentation. In this case, the attacker does not need to know the precise point of the click of the user—no matter which point in the region is clicked by the user, the object cannot be accurately segmented by SAM. In addition, we investigate to improve the transferability of the attack such that it can operate under a black-box setting. As shown in Fig. 1, after applying the adversarial perturbation to the image, clicking on any point in the yellow box (attacker-specified region) will no longer separate the corgi from the rest of the image.
@@ 3 Preliminary
$x^{(t+1)}=\textit{Clip}_{x,\epsilon}\left(x^{(t)}+\alpha\cdot\textit{sign}(\nabla_{x}L(\theta,x^{(t)},y))\right)$ (3)
where $x^{(t)}$ is the adversarial example at iteration $t$, $\alpha$ is the step size, $L$ is the loss function defined by the specific task, $\theta$ represents the model parameters, and $y$ is the ground-truth label. The function $\textit{Clip}_{x,\epsilon}$ ensures that the perturbed image remains within an $\epsilon$-neighborhood of the original image, which means $\lVert\delta\rVert_{\infty}<\epsilon$. In other words, $\epsilon$ controls the magnitude of the adversarial perturbation. The method is untargeted (i.e., the adversarial example aims to obtain any other labels that are not the ground-truth label $y$). It serves as a fundamental building block of our attack method.
@@ 4.1 Threat Movel and Attack Overview
$L(x,p)$ $=\left\|\textit{Clip}\left(\textit{SAM}(p,x+\delta;\theta),\right.\right.$
$\quad\left.\left.\textit{min}=Neg_{th}\right)-Neg_{th}\right\|^{2},$ (4)
@@
where $\delta^{*}$ represents the optimal value of $\delta$, which is the perturbation that achieves the best adversarial effect. The application of the clip function is strategic, preventing the predictive values from becoming excessively negative, which could inadvertently impede the optimization process. This ensures that predicted values, $\textit{SAM}(p,x+\delta)$, are coerced towards being less than or equal to $Neg_{th}$, with clipping applied to maintain values above this threshold.